OpenStudy (anonymous):

5log5 with an exponent of 10 how do you enter this on the TI-84 Plus?

OpenStudy (saifoo.khan):

i knw about mine.

OpenStudy (anonymous):

just as it is written. that's the beauty of the Ti-83, you can enter without using multiplication signs alot of the time.

OpenStudy (anonymous):

does that apply to the TI-48?

OpenStudy (anonymous):

make sure you close off the parenthesis behind the log 5 before you hit the ^ key.

OpenStudy (anonymous):

84 I meant

OpenStudy (anonymous):

Yes, the ti-83 and ti-84 are very similar in how they operate.

OpenStudy (anonymous):

i tried and I cannot get the right answer

OpenStudy (anonymous):

OK, try this, Hit 5, then the log button, then 5 again, then ) then ^ then 10

OpenStudy (anonymous):

.1391731292 , but the answer is 10

OpenStudy (anonymous):

Alright, it's the way you asked your question. Is the question [5 log(5)]^10 ? That is different to 5 log(5)^10. Using PEMDAS or order of operations means that the exponent gets done first. You say the answer is 10?

OpenStudy (anonymous):

umm, there is a 5 then log and then a little 5 at the bottom and then the exponent of 10. and 10 is the answer.

OpenStudy (anonymous):

Oh, crikey, now I see.............. LOL. does it look like \[5\log _{5}10\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

That was funny:)

OpenStudy (anonymous):

Alright, that's a different story. Now I need to crank out some log rules. OK. Lemme think. Your calculator doesn't have base 5. it has base 10 and base "e". Now to get around this you need to convert the log base5 of 10 to something your calc can do.

OpenStudy (anonymous):

Sorry brain fart...........

OpenStudy (anonymous):

your humor makes this more fun. The number that it could use would be....:)?

OpenStudy (anonymous):

I'm trying to find a good link for the rules for you. Give me a few ticks............(not the bloodsucking types please)

OpenStudy (anonymous):

Ha ha ha! U got it:)

OpenStudy (anonymous):

is it is the same concept as b\[\log ^{bx} \]=x?

OpenStudy (anonymous):

Just a quick search turned up this page: I'm sure there other tons of others out there that are just as great if not better. http://www.themathpage.com/aprecalc/logarithms.htm The idea is that if you have a \[5\log _{5}\] somewhere in your expression, it cancels itself out to a "1". So all you're left with is the 10 which is the answer.

OpenStudy (anonymous):

Same if you have a \[3\log _{3}14\] your answer is 14.

OpenStudy (anonymous):

Wow! u r smart! I will look at that:) can u help with this one? \[\log_{b}^{2xy} \]

OpenStudy (anonymous):

2xy may not be an exponent

OpenStudy (anonymous):

Like \[\log _{b}2xy\] ? Do they want you to expand it?

OpenStudy (anonymous):

Oh, we found a rule that applies to it! Maybe we can not bother you and do it from the book. Thank you so much for your help!! I hope you are blessed for it.

OpenStudy (anonymous):

Oh, no worries, really no bother!!

OpenStudy (anonymous):

Thank you so much! GOD bless<><